# Skagit Fisheries Enhancement Group   UMass Amherst

### v1

@@ 6.3 Ablation Studies
$L_{distill}$ Self-distillation. In LABEL:tab:distill_ablate we analyze the effects of our proposed multi-task soft distillation approach (see Sec. 4). Note that for these experiments, the starting model is ALDI++ rather than the simple model in Tab. 2. We compare our approach with the “hard” pseudo-label approach used by prior work, where teacher predictions are post-processed with non-maximum suppression and a hard confidence threshold of 0.8 [26, 34, 12, 37]. For our proposed “soft” distillation method, we first sharpen teacher predictions at both detector stages using a sigmoid for objectness predictions and a softmax for classification predictions, both with a default temperature of 1. We see that our proposed soft targets improve performance compared to hard targets.
[…]
7 Discussion and Conclusions
In this work we proposed the ALDI framework and an improved DAOD benchmarking methodology, providing a critical reset for the DAOD research community; a new dataset CFC-DAOD, increasing the diversity and real-world applicability of DAOD benchmarks; and a new method ALDI++ that advances the state-of-the-art. We conclude with key findings.
DAOD techniques are helpful, but do not achieve oracle-level performance as previously claimed [34, 8, 5]. Top-performing DAOD methods, including ALDI++, demonstrate improvements over source-only models (see Fig. 1 and Fig. 5). However, in contrast to previously-published results, no DAOD method reaches oracle-level performance, suggesting there is still room for improvement. The gap between DAOD methods and oracles is even larger for stronger architectures like VitDet. This is a promising area for future research.
A lack of transparent comparisons has incentivized incremental progress in DAOD. Most highly-performant prior works in DAOD are some combination of DANN [17] (2016) and Mean Teacher [50] (2017) plus custom training techniques. Without fair comparisons it has been possible to propose near-duplicate methods that still achieve “state-of-the-art” performance due to hyperparameter tweaks. Our method ALDI++ establishes a strong point of comparison for Align and Distill-based approaches that will require algorithmic innovation to surpass.
Validation is the elephant in the room. All of our experiments, and all previously published work in DAOD, utilize a target-domain validation set to perform model and hyperparameter selection. This violates a key assumption in unsupervised domain adaptation: that no target-domain labels are available to begin with. Prior work has shown that it may not be possible to achieve performance improvements in domain adaptation at all under realistic validation conditions [38, 39, 31]. Therefore our results (as well as previously-published work) can really only be seen as an upper bound on DAOD performance. While this is valuable, further research is needed to develop effective unsupervised validation procedures for DAOD.
Acknowledgements. This material is based upon work supported by: NSF CISE Graduate Fellowships Grant #2313998, MIT EECS department fellowship #4000184939, MIT J-WAFS seed grant #2040131, and Caltech Resnick Sustainability Institute Impact Grant “Continuous, accurate and cost-effective counting of migrating salmon for conservation and fishery management in the Pacific Northwest.” Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NSF, MIT, J-WAFS, Caltech, or RSI. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources [45]. We also thank the Alaska Department of Fish and Game for their ongoing collaboration and for providing data, and Sam Heinrich, Neha Hulkund, Kai Van Brunt, and Rangel Daroya for helpful feedback.

### v3

@@ Appendix E The ALDI Codebase
E.1 Detection Framework
We designed the ALDI codebase to be lightweight and extensible. For this reason, we build on top of a recent version of Detectron2 (Wu et al., 2019). The last tagged release of Detectron2 was v0.6 in November 2021, however there have been a number upgrades since then leading to state-of-the-art performance. Thus, we use a fixed version that we call v0.7ish based off of an unofficial pull request for v0.7, commit 7755101 dated August 30 2023. We include this version of Detectron2 as a pip-installable submodule in the ALDI codebase for now, noting that once the official version is released it will no longer need to be a submodule (i.e. it will be able to be directly installed through pip without cloning any code).
Our codebase makes no modifications to the underlying Detectron2 code, making it a lightweight standalone framework. This is in contrast to existing DAOD codebases (see Table 11) that often duplicate and modify the underlying framework as part of their implementation. By building on top of Detectron2 rather than within it, our codebase is up to 13x smaller than other DAOD codebases while providing more functionality. We note that in Table 11, other codebases implement a single method while ours supports all methods studied.
E.2 Speedups
We found significant bottlenecks in training in other Detectron2-based codebases. Notably, we found that dataloaders and transform implementations were inefficient. These included, for instance:
  • Converting tensors back and forth between torch, numpy, and PIL during augmentation. We addressed this, reimplementing transforms as needed so that everything stays in torch.
  • Using the random hue transform from torchvision. We found minimal changes in performance from disabling this component of the ColorJitter transform.
  • Using separate dataloaders for weakly and strongly augmented imagery. We instead use a single dataloader per domain, with a hook to retrieve weakly augmented imagery before strong augmentations are performed.
We reimplemented the dataloaders and augmentation strategies used by AT, MIC, and others to be more efficient, leading to a 5x speedup in training time per image compared to AT.
Appendix F Tabular Results
Table 12: Results with ResNet50-FPN backbones and ImageNet pre-training. Previously-published results are shown in gray. Best results for each benchmark are in bold and second-best are underlined.
Method CS $\rightarrow$ Foggy CS Sim10k $\rightarrow$ CS CFC Kenai $\rightarrow$ Channel
Source-only (23.5) 51.9 (35.5) 61.4 67.3
UMT (Deng et al., 2021) (41.7) 56.7 (43.1) 41.8 64.9
SADA (Chen et al., 2021) (44.0) 42.9 55.7 61.6
PT (Chen et al., 2022) (47.1) 51.5 (55.1) 64.3 67.0
MIC (Hoyer et al., 2023) (47.6) 54.0 64.3 71.9
AT (Li et al., 2022b) (50.9) 61.0 56.7 69.2
ALDI++ (Ours) 63.9 69.1 72.6
Oracle (42.7) 62.4 (66.4) 83.3 77.1
Table 13: Results with ResNet50-FPN backbones and COCO pre-training. Best results for each benchmark are in bold and second-best are underlined.
Method CS $\rightarrow$ Foggy CS Sim10k $\rightarrow$ CS CFC Kenai $\rightarrow$ Channel
Source-only 59.1 76.8 66.7
UMT (Deng et al., 2021) 61.4 58.7 61.2
SADA (Chen et al., 2021) 54.2 71.8 58.9
PT (Chen et al., 2022) 59.2 70.6 69.0
MIC (Hoyer et al., 2023) 61.7 73.1 75.5
AT (Li et al., 2022b) 63.3 72.0 69.1
ALDI++ (Ours) 66.8 77.8 76.1
Oracle 67.2 86.1 73.8
Table 14: Results with VitDet-B backbones and ImageNet pre-training. Best results are in bold and second-best are underlined.
Method CS $\rightarrow$ Foggy CS
Source-only 68.1
UMT (Deng et al., 2021) 63.9
SADA (Chen et al., 2021) 54.4
PT (Chen et al., 2022) 65.8
MIC (Hoyer et al., 2023) 70.6
AT (Li et al., 2022b) 66.6
ALDI++ (Ours) 71.0
Oracle 74.5
Appendix G Qualitative Results
We visualize predictions from all models on Cityscapes $\rightarrow$ Foggy Cityscapes in Fig. 13, Sim10k $\rightarrow$ Cityscapes in Fig. 14, and CFC-DAOD in Fig. 15.
Cityscapes $\rightarrow$ Foggy Cityscapes We choose 1 random frame from each city in the validation set (Munster, Lindau, and Frankfurt) and show the 0.005 and 0.02 fog levels. We see that, compared to ALDI++, UMT and PT consistently suffer from more false positive detections, while other methods display both false positive and false negative detections. AT and MIC perform qualitatively similary to ALDI++ in many cases. AT suffers from more false positives than ALDI++ in all locations but fewer false negatives in Lindau. MIC performs better than ALDI++ in this randomly-selected Lindau frame. In the Frankfurt example, we see one downside of the Foggy Cityscapes benchmark: since detection annotations are generated programatically from segmentations, strange false positives can exist in the ground truth. Our new dataset CFC-DAOD addresses this problem by focusing directly on object detection.
Sim10k $\rightarrow$ Cityscapes In Fig. 14 we compare results of all methods on the Sim10k $\rightarrow$ Cityscapes benchmark on two random images from each location in the Cityscapes validation set. In Munster, we see that all methods, including ALDI++, struggle with differentiating overlapping cars in the first image. In the second image, ALDI and MIC outperform other methods in differentiating cars in groups with less overlap than the first example. In Lindau, most prior work exhibits significantly more false negatives than ALDI++ and AT. AT, however, merges detections of multiple cars, leading to false negatives as well. In Frankfurt, most methods again exhibit significant false negatives, while ALDI++ and MIC show small false positive detections. Across all locations, MIC and SADA predict large false positives caused by foreground street lines.
CFC-DAOD In Fig. 15 we compare results of all methods on the CFC-DAOD Kenai $\rightarrow$ Channel benchmark. We select two random images from each camera view in the test set. These camera views represent different range windows of the sonar camera, with Stratum1 the nearest-range (leading to the highest resolution and largest fish), followed by Stratum2, and finally Stratum3 is the longest-range (lowest resolution and smallest fish). We can see that each Stratum exhibits its own challenges in differentiating fish from the background, dealing with low signal-to-noise ratios, and differences in target size. In Stratum1, all methods except SADA detect the easily identifiable large fish in the first frame, but all methods suffer from the same false negative in the second frame. ALDI++ and UMT predict the same false positive near the edge of the field of view. In Stratum2, all methods suffer from false positives caused by background texture, and UMT, SADA, and PT suffer false negatives. In Stratum3, all methods fail to detect the furthest-range fish which are highly occluded by background texture, but exhibit fewer false positives than the other strata.
Figure 13: Qualitative results on Foggy Cityscapes. Best viewed maginified.
Figure 14: Qualitative results on Sim10k $\rightarrow$ Cityscapes. Best viewed maginified.
Figure 15: Qualitative results on the CFC-DAOD test set. Best viewed maginified.